### v1

@@ 3.2.4 The disk-corona structure of AGN from X-ray polarimetry
IXPE has significantly advanced our understanding of the AGN central engine by providing the first X-ray polarization measurements of hot coronae in unobscured Seyferts [210]. These observations have helped constrain theoretical models of X-ray emission in such extreme environments. Indeed, the X-ray polarization signal strongly depends on the morphology of the corona. Radially extended configurations, such as a slab sandwiching the disk or a wedge structure, produce polarization parallel to the symmetry axis, with a degree increasing with inclination [211]. Vertically extended coronae instead produce polarization perpendicular to the symmetry axis, generally with a lower degree. IXPE results on MCG-5-23-16 [212, 213], IC 4329A [214], and NGC 4151 [180, 215] are consistent with the X-ray corona lying along the equatorial plane; however, only one solid detection ($>3\sigma$) is obtained in the case of NGC 4151.
@@ 3.2.5 Probing strong gravity effects with QPEs and TDEs
X-ray quasi-periodic eruptions (QPEs) are intensive, recurring soft X-ray bursts from galaxy nuclei. The first tentative QPE source RX J1301.9+2747 was reported in 2013 (Sun et al.), and the first confirmed QPE source was observed in 2019 in GSN 069 with nine-hour cycles of intense flares (Miniutti et al. 2019). This discovery was then followed by more detections in RX J1301.9+2747 (20-hour cycles), eRO-QPE1, eRO-QPE2, and AT2019qiz (48-hour cycles post-TDE) (Giustini et al. 2020; Arcodia et al. 2021; Nicholl et al. 2024). QPEs are preferentially found in dwarf galaxies which host SMBHs of $10^{5-7}M_{\odot}$. The sample, though small, shows diverse recurrence patterns, from regular to erratic (Quintin et al. 2024).
@@
Observationally, QPEs are defined by their dramatic X-ray variability. The eruptions are soft, with spectra peaking below 2 keV, often modeled as thermal emission from a disk with temperatures around 50 eV during quiescent phases (Miniutti et al. 2019). Burst durations range from tens of minutes to a few hours, with peak times and durations shortening at higher energies, indicating complex energy-dependent evolution (Arcodia et al. 2022). Recurrence times vary widely, from GSN 069’s nine hours to AT2019qiz’s two days, and some sources like eRO-QPE1 display irregular intervals, challenging simple periodicity models (Chen et al. 2024). Notably, some QPE hosts lack optical signatures of active galactic nuclei (AGN), suggesting these events can occur without pre-existing accretion activity (Arcodia et al. 2021).
@@
Among the diverse phenomena found in QPE observations, two illuminating discoveries reported in 2024 are particularly informative for understanding the origin of QPEs. The first discovery is the direct detection of QPEs in light curves of two recent TDEs AT 2019qiz and AT 2019vcb (Nicholl et al. 2024 and Bykov et al. 2024), confirming that at least some of QPEs are connected to TDEs. The second discovery is about the alternating long-short pattern $T_{\rm long}\ \&\ T_{\rm short}$ in the recurrence times of QPEs, which have been noticed in several QPE sources sources including the most famous GSN 069 (Miniutti et al. 2019; Giustini et al. 2020; Arcodia et al. 2021, 2022). In a recent work, Zhou et al. (2024a) further pointed out the recurrence times of GSN 069, both $T_{\rm long}$ and $T_{\rm short}$, show clear variations, while the sum of two consecutive recurrence times $T_{\rm long}+T_{\rm short}$ remains nearly a constant (see Fig. 21). This second discovery indicates that the fundamental period of QPEs is actually $T_{\rm long}+T_{\rm short}$ instead of either $T_{\rm long}$ or $T_{\rm short}$ , and two bursts are produced per period.
@@
Figure 21: Light curves from 4 observations of GSN 069 QPEs during Dec. 2018-May 2019 (Miniutti et al. 2019), where the alternating long-short pattern in the recurrence times, and the clear variations in both $T_{\rm long}$ and $T_{\rm short}$ along with the approximately constant $T_{\rm long}+T_{\rm short}$ are exhibited (Zhou et al. 2024a).
@@
Both the TDE-QPE connection and the pattern of constant $T_{\rm long}+T_{\rm short}$ favor the EMRI (extreme mass ratio inspiral)+disk model, in which a stellar object orbiting around the SMBH, colliding with the accretion disk to trigger quasi-periodic bursts (Linial et al. 2023; Franchini et al. 2023; Tagawa et al. 2023). Though alternative models, including disk instabilities and repeating partial TDEs have also been proposed as possible origins of QPEs, they are incompatible with either the TDE-QPE connection or the constant $T_{\rm long}+T_{\rm short}$ pattern in the reccurrence times.
@@
In the framework of EMRI+disk model, the variations of $T_{\rm long}$ and $T_{\rm short}$ are the natural consequences of EMRI apsidal precession in the curved spacetime. Therefore, two intrinsic periods of the EMRI trajectory are encoded in the QPE timing, the EMRI orbital period $T_{\rm obt}(\approx T_{\rm long}+T_{\rm short})$ and the EMRI apsidal precession period $T_{\rm prec}$ in which the recurrence times are modulated. Several previous works have attempted to fit the QPE timing data with EMRI geodesics (Xian et al. 2021, Franchini et al. 2023, Zhou et al. 2024a,b) or EMRI trajectories that are perturbed by the accretion disk (Zhou et al. 2024c). Taking GSN 069 as an example, Zhou et al. (2024c) found an orbital period $T_{\rm obt}\approx 65$ ks, an apsidal precession period $T_{\rm prec}\approx 76$ days, and an orbital eccentricity $e\approx 0.04$. The two periods are related to the EMRI orbital parameters as
$T_{\rm obt}=2\pi\left(\frac{A}{M_{\bullet}}\right)^{3/2}M_{\bullet}\ ,\quad\frac{T_{\rm prec}}{T_{\rm obt}}=\frac{p}{3M_{\bullet}}=\frac{A}{3M_{\bullet}}(1-e^{2})\ ,$ (3)
@@
where geometric units with $G=c=1$ are used, $M_{\bullet}$ is the gravitational radius of the SMBH, $A$ is the semi-major axis and $p$ is the semi-latus rectum. From the two periods inferred from QPE timing data, it is straightforward to obtain the value of the SMBH mass $M_{\bullet}$. This QPE timing method is promising in accurate measurement of SMBH masses (Zhou et al. 2024c). Additionally, stellar transits across a magnetized accretion flow that quasi-periodically ejects plasmoids has also been put forward as a promising mechanism [216].
@@
Polarization studies for QPEs are under-explored, but theoretical models offer some predictions. In the stellar interaction model, if eruptions arise from localized disk collisions, the X-ray emission could be polarized due to scattering in an asymmetric geometry, such as a hot spot or outflow. The degree of polarization would depend on the viewing angle and disk structure, potentially reaching a few percent in soft X-rays, as seen in some accretion disk simulations (Schnittman et al. 2020). For disk instability models, polarization might be lower or absent unless the instability triggers anisotropic outflows or a corona, which could scatter photons and induce polarization (Marinucci et al. 2021). No direct polarization measurements exist yet for QPEs, as current X-ray polarimeters like IXPE lack the sensitivity for such short and faint events.
Based on eXTP’s expected orbital parameters and instrument performance, we conducted simulations for the detectability of typical QPEs by eXTP. ???
@@
A tidal disruption event (TDE) occurs when a star ventures within the tidal radius of a supermassive black hole (SMBH; [217, 218]. These events generate flares lasting months to years, exhibiting extreme variability across multi-wavelengths [219] and revealing previously dormant SMBHs. For approximately a year after the disruption, the accretion rate remains high, resulting in a geometrically thick nascent accretion disk. Due to the typical misalignment between the disrupted star’s orbit and the BH spin axis, strong viscous forces and frame-dragging effects cause the entire disk to undergo rigid-body Lense-Thirring precession around the BH spin axis (e.g., [220, 221]). For an accretion disk extending from the ISCO to approximately the tidal disruption radius, the precession period varies from one to tens of days, depending on the SMBH mass and spin parameter ([222, 223]). In the case of intermediate-mass black holes, the precession period can be significantly shorter, e.g. 85 s in 3XMM J215022.4–055108 [224].
@@
Unlike in AGNs and XRBs, coronae are typically absent in TDEs, at least during the early stages post-disruption, making disk reflection undetectable in most cases. As a result, periodic X-ray variations provide the most viable method for measuring the BH spin parameter. This approach has been applied to TDEs such as Swift J1644 ([222, 225] and AT2020ocn [226], which exhibited periodicities of 2.7 days and 15 days, respectively, attributed to jet and disk precession. GRMHD simulations predict that both alignment and precession should gradually slow down over time due to angular momentum transport and radial disk expansion [227]. However, these predictions remain unverified, particularly at late evolutionary stages of TDEs, due to the limited statistical constraints from current instruments. The large effective area of eXTP/SFA in the soft X-ray band makes it an ideal instrument to thoroughly investigate the evolution of X-ray periodicity and its associated spectral changes. More precise measurements on these quantities could further test the Bardeen–Petterson effect [181], which dampens LT precession, thereby providing key constraints on disk dynamics as it transitions from geometrically thick to geometrically thin.
@@
When the stellar orbital periastron $R_{p}$ is very close to the BH, i.e., < $<\leavevmode\nobreak\ 6\,R_{g}$, the strong GR effect causes the orbital dynamics of the stellar debris dramatically different from that in ordinary TDEs, thus rendering the term extremely relativistic TDEs (eTDEs). Such regime dominates those TDEs with BH mass in the range of $10^{7}-10^{8}\leavevmode\nobreak\ M_{\odot}$. Particularly, the strong apsidal precession is so strong that the debris will wind a few times at $R_{p}$ around the BH, before its path extends out to large distances (see Figure 22). Internal shocks occur within the wound debris stream, dissipating energy, and followed by the accretion of the compact, circularized flow, as was shown by simulation in [228].
Figure 22: The path of the stellar debris in a typical extremely relativistic TDE (eTDE). The solid white curve depicts the geodesic of an orbit with $R_{p}=4.03\,R_{g}$ around an SMBH (red disk at the center); the arrow indicates the orbital direction. Figure from [228].
@@ 4 Summary
The study of accreting black holes provides insights into some of the most fundamental objects and processes that impact our Universe, while presenting great theoretical and observational challenges. Black holes themselves are a manifestation of gravity at its most extreme and provide a setting which allows us to explore our best theory of gravity - General Relativity - and test the spacetime metric. The physics of accretion and the outflows it produces lead us to couple our understanding of electromagnetism and plasma physics, thermodynamics and radiative processes to this deeper theory. This synthesis allows us to use electromagnetic observations as probes of the behaviour of matter accreting in strong gravity, and gravity itself, but is incredibly demanding in terms of the quality of data required.
@@
    The combination of collecting area, sensitivity, spectral-response and high orbit of eXTP-SFA make it ideal for the detection and study of quasi-periodic signals from supermassive black holes (SMBH) which can result from a variety of physical phenomena (Section 3.2). Dynamical-time-scale QPOs can be easily detected from even faint AGN. Strong constraints on orbital dynamics of hotspots orbiting the Sgr A SMBH in our own galaxy can be obtained by their X-ray flare profiles. The soft emission from accretion disks produced by tidal disruption of stars can be probed for precession signals and the QPEs now associated with repeated disruption of the disk by the stellar orbits.
@@ Acknowledgements.
AV acknowledges support from the Academy of Finland grant 355672. Nordita is supported in part by NordForsk. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100).

### v2

@@ 3.2.4 The disk-corona structure of AGN from X-ray polarimetry
IXPE has significantly advanced our understanding of the AGN central engine by providing the first X-ray polarization measurements of hot coronae in unobscured Seyferts [230]. These observations have helped constrain theoretical models of X-ray emission in such extreme environments. Indeed, the X-ray polarization signal strongly depends on the morphology of the corona. Radially extended configurations, such as a slab sandwiching the disk or a wedge structure, produce polarization parallel to the symmetry axis, with a degree increasing with inclination [100]. Vertically extended coronae instead produce polarization perpendicular to the symmetry axis, generally with a lower degree. IXPE results on MCG-5-23-16 [231, 232], IC 4329A [233], and NGC 4151 [200, 234] are consistent with the X-ray corona lying along the equatorial plane; however, only one solid detection ($3\sigma$) is obtained in the case of NGC 4151.
@@ 3.2.5 Probing strong gravity effects with QPEs and TDEs
X-ray quasi-periodic eruptions (QPEs) are intensive, recurring soft X-ray bursts from galaxy nuclei. The first tentative QPE source RX J1301.9+2747 was reported in 2013 [235], and the first confirmed QPE source was observed in 2019 in GSN 069 with nine-hour cycles of intense flares [236]. This discovery was then followed by more detections in RX J1301.9+2747 (20-hour cycles), eRO-QPE1, eRO-QPE2, and AT2019qiz (48-hour cycles post-TDE) [238, 239, 237]. QPEs are preferentially found in dwarf galaxies which host SMBHs of $10^{5-7}M_{\odot}$. The sample, though small, shows diverse recurrence patterns, from regular to erratic [240].
@@
Observationally, QPEs are defined by their dramatic X-ray variability. The eruptions are soft, with spectra peaking below 2 keV, often modeled as thermal emission from a disk with temperatures around 50 eV during quiescent phases [236]. Burst durations range from tens of minutes to a few hours, with peak times and durations shortening at higher energies, indicating complex energy-dependent evolution [241]. Recurrence times vary widely, from GSN 069’s nine hours to AT2019qiz’s two days, and some sources like eRO-QPE1 display irregular intervals, challenging simple periodicity models. Notably, some QPE hosts lack optical signatures of active galactic nuclei (AGN), suggesting these events can occur without pre-existing accretion activity [239].
@@
Among the diverse phenomena found in QPE observations, two illuminating discoveries reported in 2024 are particularly informative for understanding the origin of QPEs. The first discovery is the direct detection of QPEs in light curves of two recent TDEs AT 2019qiz and AT 2019vcb ([237] [242]), confirming that at least some of QPEs are connected to TDEs. The second discovery is about the alternating long-short pattern $P_{\rm long}\ \&P_{\rm short}$ in the recurrence times of QPEs, which have been noticed in several QPE sources sources including the most famous GSN 069 [236, 238, 239, 241]. In a recent work, Zhou et al. [243] further pointed out the recurrence times of GSN 069, both $P_{\rm long}$ and $P_{\rm short}$, show clear variations, while the sum of two consecutive recurrence times $P_{\rm long}+P_{\rm short}$ remains nearly a constant (see Fig. 21). This second discovery indicates that the fundamental period of QPEs is actually $P_{\rm long}+P_{\rm short}$ instead of either $P_{\rm long}$ or $P_{\rm short}$, and two bursts are produced per period.
@@
Figure 21: Light curves from 4 observations of GSN 069 QPEs during Dec. 2018-May 2019 [236], where the alternating long-short pattern in the recurrence times, and the clear variations in both $P_{\rm long}$ and $P_{\rm short}$ along with the approximately constant $P_{\rm long}+P_{\rm short}$ are exhibited [243, 251].
@@
Both the TDE-QPE connection and the pattern of constant $P_{\rm long}+P_{\rm short}$ favor the EMRI (extreme mass ratio inspiral)+disk model, in which a stellar object orbiting around the SMBH, colliding with the accretion disk to trigger quasi-periodic bursts [245, 244, 246]. Though alternative models, including disk instabilities and repeating partial TDEs have also been proposed as possible origins of QPEs, they are incompatible with either the TDE-QPE connection or the constant $P_{\rm long}+P_{\rm short}$ pattern in the recurrence times. Additionally, stellar transits across a magnetized accretion flow that quasi-periodically ejects plasmoids has also been put forward as a promising mechanism [247].
@@
In the framework of EMRI+disk model, the variations of $P_{\rm long}$ and $P_{\rm short}$ are the natural consequences of EMRI apsidal precession in the curved spacetime. Therefore, two intrinsic periods of the EMRI trajectory are encoded in the QPE timing, the EMRI orbital period $P_{\rm obt}\approx(P_{\rm long}+P_{\rm short})$ and the EMRI apsidal precession period $P_{\rm prec}$ in which the recurrence times are modulated. Several previous works have attempted to fit the QPE timing data with EMRI geodesics ([249, 244, 243, 248]) or EMRI trajectories that are perturbed by the accretion disk [250]. Taking GSN 069 as an example, Zhou et al. [250] found an orbital period $P_{\rm obt}\approx 65$ ks, an apsidal precession period $P_{\rm prec}\approx 76$ days, and an orbital eccentricity $e\approx 0.04$. The two periods are related to the EMRI orbital parameters as
$P_{\rm obt}=2\pi\left(\frac{A}{M_{\bullet}}\right)^{3/2}M_{\bullet}\ ,\quad\frac{P_{\rm prec}}{P_{\rm obt}}=\frac{p}{3M_{\bullet}}=\frac{A}{3M_{\bullet}}(1-e^{2})\ ,$ (3)
@@
where geometric units with $G=c=1$ are used, $M_{\bullet}$ is the gravitational radius of the SMBH, AA is the semi-major axis and $p$ is the semi-latus rectum. From the two periods inferred from QPE timing data, it is straightforward to obtain the value of the SMBH mass $M_{\bullet}$. This QPE timing method is promising in accurate measurement of SMBH masses [250, 251].
@@
Polarization studies for QPEs are under-explored, but theoretical models offer some predictions. In the stellar interaction model, if eruptions arise from localized disk collisions, the X-ray emission could be polarized due to scattering in an asymmetric geometry, such as a hot spot or outflow. The degree of polarization would depend on the viewing angle and disk structure, potentially reaching a few percent in soft X-rays, as seen in some accretion disk simulations [104]. For disk instability models, polarization might be lower or absent unless the instability triggers anisotropic outflows or a corona, which could scatter photons and induce polarization [231]. No direct polarization measurements exist yet for QPEs, as current X-ray polarimeters like IXPE lack the sensitivity for such short and faint events.
@@
A tidal disruption event (TDE) occurs when a star ventures within the tidal radius of a supermassive black hole [252, 253]. These events generate flares lasting months to years, exhibiting extreme variability across multi-wavelengths [254] and revealing previously dormant SMBHs. For approximately a year after the disruption, the accretion rate remains high, resulting in a geometrically thick nascent accretion disk. Due to the typical misalignment between the disrupted star’s orbit and the black hole spin axis, strong viscous forces and frame-dragging effects cause the entire disk to undergo rigid-body Lense-Thirring precession around the black hole spin axis (e.g., [255, 256]). For an accretion disk extending from the ISCO to approximately the tidal disruption radius, the precession period varies from one to tens of days, depending on the SMBH mass and spin parameter ([257, 258]). In the case of intermediate-mass black holes, the precession period can be significantly shorter, e.g. 85 s in 3XMM J215022.4–055108 [259].
@@
Unlike in AGNs and BHXRBs, coronae are typically absent in TDEs, at least during the early stages post-disruption, making disk reflection undetectable in most cases. As a result, periodic X-ray variations provide the most viable method for measuring the black hole spin parameter. This approach has been applied to TDEs such as Swift J1644 ([257, 260] and AT2020ocn [261], which exhibited periodicities of 2.7 days and 15 days, respectively, attributed to jet and disk precession. GRMHD simulations predict that both alignment and precession should gradually slow down over time due to angular momentum transport and radial disk expansion [262]. However, these predictions remain unverified, particularly at late evolutionary stages of TDEs, due to the limited statistical constraints from current instruments. The large effective area of eXTP/SFA in the soft X-ray band makes it an ideal instrument to thoroughly investigate the evolution of X-ray periodicity and its associated spectral changes. More precise measurements on these quantities could further test the Bardeen–Petterson effect [201], which dampens LT precession, thereby providing key constraints on disk dynamics as it transitions from geometrically thick to geometrically thin.
@@
When the stellar orbital periastron $R_{p}$ is very close to the black hole, i.e., < $\leavevmode\nobreak\ 6\,R_{g}$, the strong GR effect causes the orbital dynamics of the stellar debris dramatically different from that in ordinary TDEs, thus rendering the term extremely relativistic TDEs (eTDEs). Such regime dominates those TDEs with black hole mass in the range of $10^{7}-10^{8}\leavevmode\nobreak\ M_{\odot}$. Particularly, the strong apsidal precession is so strong that the debris will wind a few times at $R_{p}$ around the black hole, before its path extends out to large distances  [263]. Internal shocks occur within the wound debris stream, dissipating energy, and followed by the accretion of the compact, circularized flow, as was shown by simulation in [263].
@@ 4 Summary
The study of accreting black holes provides insights into some of the most fundamental objects and processes that impact our Universe, while presenting great theoretical and observational challenges. Black holes themselves are a manifestation of gravity at its most extreme and provide a setting which allows us to explore our best theory of gravity - GR - and test the spacetime metric. The physics of accretion and the outflows it produces lead us to couple our understanding of electromagnetism and plasma physics, thermodynamics and radiative processes to this deeper theory. This synthesis allows us to use electromagnetic observations as probes of the behaviour of matter accreting in strong gravity, and gravity itself, but is incredibly demanding in terms of the quality of data required.
@@
    The combination of collecting area, sensitivity, spectral-response and high orbit of eXTP-SFA make it ideal for the detection and study of quasi-periodic signals from supermassive black holes which can result from a variety of physical phenomena (Section 3.2). Dynamical-time-scale QPOs can be easily detected from even faint AGN. Strong constraints on orbital dynamics of hotspots orbiting the Sgr A supermassive black hole in our own galaxy can be obtained by their X-ray flare profiles. The soft emission from accretion disks produced by tidal disruption of stars can be probed for precession signals and the QPEs now associated with repeated disruption of the disk by the stellar orbits.
@@ Acknowledgements.
This work is supported by China’s Space Origins Exploration Program. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). AV acknowledges support from the Academy of Finland grant 355672. Nordita is supported in part by NordForsk.
Conflict of Interest
The authors declare that they have no conflict of interest.